Using the quadratic equation y = x2 - 6x + 8 = 0, perform the following tasks:
Solve by factoring.
Solve by using the quadratic formula.

Answer by ptaylor(1894)

Using the quadratic equation y = x2 - 6x + 8 = 0, perform the following tasks:
(1) Solve by factoring.
(2) Solve by using the quadratic formula.

(1) This quadratic equation is in standard form (ax^2+bx+c=0) with:
a=1
b=-6
c=8
Now when a=1, then b is the sum of the factors of c, if the quadratic is factorable. What are the factors of 8 that will add up to -6?????
Well, it's (-4)(-2)
So x^2 - 6x + 8=0 equals
(x-4)(x-2)=0 you can multiply this out to verify
x=4
x=2
(2)
Quadratic formula: x=(-b+or-sqrt(b^2-4ac))/2a
x=(6+or-sqrt(36-32))/2
x=(6+or-sqrt4)/2
x=-(6+or-2)/2 divide top and bottom by 2
x=3+or-1
x=3+1=4
x=3-1=2

Hope this helps---ptaylor